# CLIC: Contrastive Learning Framework for Unsupervised Image Complexity Representation

### v1

@@ E Future Work
This paper proposes CLIC, an unsupervised image complexity representation learning framework based on comparative learning. Unlike previous approaches that rely on well-labeled datasets, CLIC can efficiently learn IC features based on unlabeled data. Through extensive experimental validation, the results show that CLIC can learn effective IC representations and performs competitively with supervised learning methods for fine-tuning on the IC9600. In addition, CLIC shows significant improvement in several downstream tasks, demonstrating its potential in real-world scenarios.
Although CLIC can efficiently learn image complexity representations, we believe it is flawed.
@@
  • First, we introduce image prior information as a supervised signal to construct an auxiliary task that helps CLIC learn image complexity representation. However, the contrast loss is still based on InfoNCE, which may inevitably cause CLIC to learn a few categories and object attributes.
@@
  • Second, the trained CLIC needs to parallelize the query encoder with them (backbones) when applying to downstream tasks. This can lead to a loss of efficiency in downstream tasks.
@@
Thus our future work has two parts. We will build a pretext task suitable for image complexity representation learning, which will be more coupled than normal contrast loss. On the other hand, we are going to build a framework for image complexity knowledge distillation, which is expected to migrate the IC knowledge of the query encoder to the downstream task model. This will keep the downstream task efficient while increasing the performance

### v2

@@ 6 Limitations
This paper proposes CLIC, an unsupervised image complexity representation learning framework based on comparative learning. Unlike previous approaches that rely on well-labeled datasets, CLIC can efficiently learn image complexity (IC) features from unlabeled data. Through extensive experimental validation, we demonstrate that CLIC can effectively learn IC representations and performs competitively with supervised learning methods when fine-tuned on the IC9600 dataset. Furthermore, CLIC significantly improves performance in several downstream tasks, showcasing its potential for real-world applications.
However, we acknowledge a few limitations of CLIC:
  • 1.
@@
    First, while we introduce image prior information as a supervised signal to construct an auxiliary task that aids CLIC in learning image complexity, the contrastive loss still relies on InfoNCE. This may lead CLIC to inadvertently focus on certain categories or object attributes, which could hinder the pure learning of image complexity.
  • 2.
@@
    Second, when applying CLIC to downstream tasks, the query encoder needs to be parallelized with the backbone models. This requirement can result in a loss of efficiency during downstream task execution.
@@
For future work, we have two main directions. First, we aim to develop a more suitable pretext task for image complexity representation learning, which will be more tightly coupled than the traditional contrastive loss. Second, we plan to establish a framework for image complexity knowledge distillation, which would allow the IC knowledge from the query encoder to be transferred to downstream task models. This approach will maintain the efficiency of downstream tasks while enhancing their performance.
7 Conclusion
Image complexity assessment plays a crucial role in optimizing computer vision tasks. This paper introduces CLIC, an unsupervised image complexity representation learning framework based on comparative learning. Unlike traditional approaches that depend on well-labeled datasets, CLIC efficiently learns image complexity (IC) features from unlabeled data. Extensive experimental validation demonstrates that CLIC can effectively learn IC representations, performing competitively with supervised learning methods when fine-tuned on IC9600. Furthermore, CLIC shows substantial improvements in several downstream tasks, highlighting its potential for real-world applications. Looking ahead, our future work will focus on developing more suitable pretext tasks for IC representation learning, ensuring that these tasks remain free from interference by attributes such as categories or objects. Additionally, we aim to create an image complexity knowledge transfering framework that could serve as a ”free lunch” for optimizing downstream tasks, further improving efficiency and performance.
8 Acknowledgements
This work would not have been possible without financial support from the National Natural Science Foundation of China (No.51209167; 52178393; 51578447), Science and Technology Innovation Team of Shaanxi Innovation Capability Support Plan (2020TD005), Xi’an Scientists & Engineers Workforce Building Project (2024JH-KGDW-0112). The authors are also grateful to the editors and anonymous reviewers for their sound and valuable suggestions on the manuscript.
Appendix A Datasets
A.1 ImageNet
A.2 Flickr
Flickr [43] contains $\sim$ 5 billion images of multiple types from Flickr (a popular online photo and video hosting, sharing, and management platform), covering a rich diversity of scenes and content. These images range from common scene types, such as natural scenes from everyday life, portraits, buildings, and city streetscapes, to more creatively expressive content, such as still life and artistic photography. The complexity of the images shows significant differences in the number of visual elements, distribution, texture, light and shadow changes, etc., which is suitable for image complexity analysis and evaluation.
A.3 IC9600
Figure 1: CLIC fint-tuning pipeline.
Appendix B Metrics Details
B.1 Unsupervised Activation Energy
B.2 Edge Density
B.3 Compress Ratio
$CR=\frac{B}{H(I)}\quad\\ \quad H(I)=-\sum_{x=0}^{255}p(x)\log_{2}p(x)$ (10)
Appendix C Fine-tuning pipeline
The fine-tuning pipeline of CLIC for image complexity evaluation is in Fig.1. ICNet contains two ResNet branch (R18) to capture detail ($512\times 512\times 3$) and context ($256\times 256\times 3$) features, respectively. Followed by IC map head and IC fraction head, which are used to calculate the loss. CLIC fine-tuning only uses the query encoder (Q Enc) as a separate branch and adopts an input size of $224\times 224\times 3$, followed by the structure is same as ICNet.